madamerockles:

Factor: 36a^2b^100-1

Elsa213:

Welcome to QuestionCove. What is your question? o:

madamerockles:

i am stuck on this one question that wants me to factor 36a^2b^100-1

Hero:

Elsa, she already posted her problem.

Hero:

@madamerockles are you familiar with difference of squares formula?

madamerockles:

not really i missed alot of school so im really confuse on most of my work

Hero:

The difference of squares formula says \(a^2 - b^2 = (a + b)(a - b)\) In this case we can re-write the given problem in the following manner: \(36a^2 b^{100} - 1 = (6ab^{10})^2 - 1^2\)

madamerockles:

oh okk i understand now thank you

Hero:

In other words, in this case: let \(a = 6ab^{10}\) and \(b = 1\) Thus you should factor and get: \((6ab^{10} - 1)(6ab^{10}+1)\)